# BINDy – Bayesian identification of nonlinear dynamics with reversible-jump Markov-chain Monte-Carlo

### v1

@@ 3.3 Bayesian identification of nonlinear dynamics
$a=a^{(0)}+\frac{n}{2}$ (24)
$b=b^{(0)}+\frac{\sum_{i}^{n}(\dot{x}-\Theta(x)\Xi)^{2}}{2}$ (25)
where $n$ is the number of data in $\dot{x}$. The overall sampler is summarised in Algorithm 1.
Algorithm 1 Gibbs sampler with reversible-jump model moves
@@
Parameter prior mean and variance $\mu_{m}^{(0)},\Sigma_{m}^{(0)}$, variance prior parameters $a^{(0)},b^{(0)}$, prior over models $p(m)$, jump transition probabilities $J(m^{\prime}|m)$.
Initialise the chain at $m=\mathcal{M}$ (i.e. with every term included) and $\sigma_{0}^{2}$
@@
     Propose a model move $m\rightarrow m^{\prime}$ by sampling from $J(m^{\prime}|m)$
     Accept the model move with probability $\alpha(\Xi\rightarrow\Xi^{\prime})$ given by (21) else remain in place.
     Propose a new noise variance $\sigma^{2}$ using the Gibbs update given by (23)
@@
return samples from $p(\Xi,m|\mathcal{D})$
@@ 3.4 Related work
The method presented herein is also not the first use of RJMCMC in the Bayesian identification of dynamic models. Troughton and Godsill [22] and Dahlin et al.  [29] consider an RJMCMC scheme for the identification of linear autoregressive models, and a RJMCMC scheme is employed in [30] to select between two candidate nonlinear models. However, to the author’s best knowledge, this work represents the first application of RJMCMC to the SINDy-type formulation of [1], and the first to explicitly enumerate a full posterior distribution over both model terms and their parameterisation jointly.
@@ 4 Results
In all case-study examples, the prior structure of the inference is set up to be weakly uninformative. In particular, the values,
$\mu_{m}\triangleq\bm{0}^{(0)},\quad\Sigma_{m}^{(0)}\triangleq 10^{3}\times I$ (26)
$a^{(0)}\triangleq 0,\quad b^{(0)}\triangleq 0$ (27)
are chosen. The prior distribution over models is selected initially to be flat with $p(m)\propto 1\ \forall\ m\in\mathcal{M}$. The authors would note that this choice not promote model parsimony a priori. For the jump transitions $J(m^{\prime}|m)$ the authors propose a ‘bit-flipping’ scheme whereby in each proposal step a random index in $[1,n]$ (where $n$ is the number of rows in $\Theta$) is selected with equal probability. If that term is present in the current model then it is removed, if not then it is added. If a move is rejected then a new index is drawn, again with even probability. The chain is initialised with all terms in $\Theta$ present in the model and with the noise variance set to $\sigma^{2}=1$. The Gibbs sampler is run for $6\times 10^{3}$ steps in total, with the first thousand samples discarded to remove the effects of any transient ‘burn-in’ behavior. The authors remark that this length of chain appears more than adequate in terms of convergence to the target distribution.
@@ 4.1 Legendre polynomials
The first case-study is a static sparse polynomial regression problem in which $\Theta(x)$ is static and is set to be the first 10 Legendre polynomials on the interval $x\in[-1,1]$. In order to demonstrate the recovery of small-valued parameters, a coefficient vector is selected with both zero and near-zero elements. The coefficient vector used here is given by,
$\Xi^{*}\triangleq\left[0.549,0,0.603,0.545,0.424,0.006,0,0,0,0.004\right]^{\mkern-1.5mu\mathsf{T}}$ (28)
@@
$y\triangleq\Theta(x)\Xi^{*}+\epsilon\quad\epsilon\sim\mathcal{N}(\bm{0},\sigma^{2}_{n}I)$ (29)
@@
The performance of both E-SINDy and the proposed method are compared in Figure 1. As can be seen in the figure, Both the proposed method and E-SINDy methods are able to correctly assign low inclusion probability to the terms in $\Xi^{*}$ equal to zero. However, the STLSQ algorithm is unable to identify terms with parameters lower than the threshold parameter $\lambda$, which in this case study is set to the default value of 0.1. While this is a simplified example, it serves to demonstrate a particular pathology, which may appear when using parameter value as a proxy for term importance. However, the E-SINDy method shows multi-modality, producing some samples in clusters away from the true value. This can be attributed to the library bagging operation. Every subset of the library functions produces a sample (none are rejected - there is no Metropolisation step). This has the effect that some samples are generated with important terms missing, biasing the parameter values of the remaining terms.
@@ 4.2 Lynx-hare population dynamics
$\begin{split}\frac{dL}{dt}&=aL+bHL\\ \frac{dH}{dt}&=cH+dHL\end{split}$ (30)
Where $a,b,c,d$ are unknown constants in the model. The dataset represents a significant challenge in that only 21 measurements are available.
@@ 4.3 Lorenz attractor dynamics
The third case study considers the chaotic dynamics of the Lorenz attractor. The dynamics of the Lorenz attractor are well studied in the SINDy literature [1, 16], and have become a benchmark system for advancements to the SINDy method. The dynamics are driven by the following system of ordinary differential equations.
$\begin{split}\frac{dx}{dt}&=10(x_{2}-x_{1})\\ \frac{dy}{dt}&=x_{1}(28-x_{3})\\ \frac{dz}{dt}&=x_{1}x_{2}-\frac{8}{3}x_{3}\end{split}$ (31)
@@
For the case study in this work, the Lorenz equations above are simulated for 10 seconds (with a further 5 seconds of unseen data for extrapolation) from an initial condition of $[x_{1},x_{2},x_{3}]^{\mkern-1.5mu\mathsf{T}}=[-8,7,27]^{\mkern-1.5mu\mathsf{T}}$ at a sample rate of $10^{2}$ Hz. Overall, $10^{3}$ points are available for training. As before, $\Theta(x)$ is set as a polynomial library, considering terms up to third order. Before training, noise at 2.5% RMS is added to the state variables. The state derivatives are computed numerically using a polynomial smoothed finite-difference scheme with a window size of 5 and a polynomial order of 3. For the SINDy results, the threshold parameter is set to 0.2 (as is done in [16]). All other parameters are unchanged from the previous case-studies.
@@
For this third case study, the effect of the prior over models $p(m)$ is also considered. In addition to the flat prior used in the previous case-study examples, a regularising geometric prior over the total number of non-zero elements in the model is considered. The probability mass function of the regularising prior is given by,
$p(m)=(1-\theta)^{d}\theta$ (32)
@@
where $d$ is the number of terms in model $m$, and $\theta$ is a hyperparameter, set to the strongly regularising value of 0.99 in this work. The effect of this prior over models is to strongly penalise (by assigning low prior probability) models with many terms.
@@
The result of the UQ for the Lorenz equations is compared in Figure 5 for BINDy (with and without the regularising prior) and E-SINDy. It can be seen in the figure that the unregularised BINDy assigns high posterior probability to many spurious terms. In contrast, Both E-SINDy and the regularised BINDy assign high confidence to terms in the Lorenz equations. This is not an unexpected result. The nature of the Lorenz equations, and the use of the polynomial library results in many, highly-correlated basis functions in $\Theta(x)$. This, plus the effect of noise permit many incorrect (but low error) parametrisations of the observed dynamics. Both E-SINDy and regularised BINDy are able to overcome this difficulty by explicitly promoting model parsimony. Because the unregularised BINDy enforces no such cognitive bias, it is unsurprising that the identified posterior contains many terms. Summary statistics of the three identified posteriors are provided in Table 1.
@@
As before, sampled model trajectories (integrated forward in time) are plotted for all three methods in Figure 6. The chaotic nature of the Lorenz equations makes tracking the true dynamics highly challenging. The well-known ‘lobe-switching’ behavior of the dynamics means that small errors can quickly cause significant deviations as the dynamics become governed by a different local attractor. This effect is clearly present in the identified distributions, which quickly become multimodal after a few oscillations. Qualitatively, it is observed that the regularised BINDy samples remain in step with the true dynamics for the longest, with E-SINDy the first to diverge. This divergence can be seen at $t<1$s where orage lines deviate from the observed data. Indeed, this is supported by the numerical results in Table 2 which show lower error statistics for the regularised BINDy approach. In the extrapolation region, the distributions have become very diffuse, in keeping with what might be expected for forecasting a chaotic attractor such as the Lorenz equations.
@@ 5 Discussion
As has been demonstrated in the third case-study of this work, the choice of prior over models $p(m)$ may be critical in enforcing parsimony. Here, a strongly regularising geometric distribution parameterised by the number of terms in the model has been applied successfully. However, vastly many more parametrisations are possible. It is expected that prior selection in practice will be driven by expert domain knowledge, wherever possible. For example, one could set the prior over each term directly, assigning high prior probability to terms that are expected to be in the model and low probabilities otherwise. Alternatively one could promote parsimony in some types of library functions while permitting dense models in other types. The flexible nature of the Bayesian formulation presented in this work, allows modellers to encode both domain knowledge and cognitive biases a priori.
@@
A related area for future investigation is the choice of the ‘jump’ kernel $J(m^{\prime}|m)$. The ‘bit-flipping’ method applied herein has proved to be effective in the case-studies considered so far, but it might be ineffective when significant multi-modality is expected in the posterior, for example when very differently parameterised models describe the data well, low probability intermediate steps might inhibit exploration of the posterior. Instead, it is interesting to imagine more sophisticated schemes that could better suit problems of this type. Also interesting are so-called ‘non-reversible’ jump methods [33] that use an adaptive scheme to ensure that the posterior is explored efficiently.

### v2

@@ 3 Bayesian identification of nonlinear dynamics
$a=a^{(0)}+\frac{n}{2}{\color[rgb]{0,0,0},}$ (25)
$b=b^{(0)}+\frac{\sum_{i}^{n}(\dot{x}-\Theta(x)\Xi)^{2}}{2}{\color[rgb]{0,0,0},}$ (26)
where $n$ is the number of data in $\dot{x}$.
3.1 A framework for Bayesian identification of nonlinear dynamics
It is useful to consider here, how the above may be used by the practitioner in order to arrive at a useful estimate of the posterior distribution $p(\Xi,m,\sigma^{2}|\dot{x},\Theta(x))$ in the context of nonlinear system identification. In practice there are several quantities that must be selected by the user. These are: the parameter and noise variance priors, the model prior, and the jump kernel. The choices of these objects will affect the cognitive biases in the inference (including the sparsity of the posterior) as well as affecting the convergence properties of the sampling scheme. Some discussion on how these objects might be selected in practice is held here.
Parameter and noise variance priors, $p(\Xi)$ and $p(\sigma^{2})$
A constraint of the method introduced here is that these priors must be conditionally conjugate to the Gaussian likelihood model in (3). Although this is a strong constraint it remains possible to reflect many types of prior belief with these objects. For the parameter priors, vague belief can be introduced by consideration of very wide Gaussian distributions centred around zero. Stronger belief can be imposed by the consideration of more concentrated variances. For the noise variance prior, lack of belief can be imposed by the improper parametrisation of the Inverse gamma distribution, $\mathcal{IG}(0,0)$ (although there are potentially limitations if samples are required to be drawn from the prior). Stronger belief can be enforced by more concentrated parameterisations.
Model space priors, $p(m)$
One of the major advantages of the formulation considered here is that prior belief can be applied directly in the model space without using parameter values as a proxy. The RJMCMC scheme employed here places no restriction on the form of the prior distribution $p(m)$. Thus, very many cognitive biases can be injected into the inference scheme by the careful consideration of this parameter. Ignorance as to the model structure and level of sparsity can be trivially introduced by the consideration of a flat model prior. For example,
${\color[rgb]{0,0,0}p(m)\propto 1,\quad\forall\ m\in\mathcal{M}.}$ (27)
The cognitive bias of model parsimony can also be introduced. For example one could naively choose to prefer models with fewer terms (thus promoting sparsity) by the consideration of a monotonically decreasing discrete density defined over the number of terms in the model. Appropriate choices might include the geometric distribution,
${\color[rgb]{0,0,0}p(m)=(1-\theta)^{d}\theta,}$ (28)
where $\theta$ is a hyperparameter that controls the extent to which more parsimonious models are preferred. Other types of belief could be applied by assigning probabilities to model term inclusion directly. An instructive example might be in the identification of structural dynamics in a near linear regime. One could envisage a prior structure that placed a high probability of selecting expected linear terms (corresponding to inertia, viscous damping and linear elasticity), while assigning less prior probability to nonlinear terms. The effect would be that the nonlinear terms would only appear in the posterior if there was significant evidence in the likelihood. As with all prior selection in Bayesian approaches, the choice of the model-space prior must ultimately be guided by the problem at hand and domain knowledge (or lack thereof) [27]. However, the authors note that in most SINDy-type applications, practitioners expect the observed dynamics to be sparse in the library of functions. This assumption motivates the use of regularising priors in general
Jump kernel, $J(m^{\prime}|m)$
The selection of the jump kernel is one of the key components of the RJMCMC and may strongly affect the convergence of the sampler. Here, we propose the following ‘bit-flipping’ scheme. In each proposal step a random index in $[1,n]$ (where $n$ is the number of rows in $\Theta$) is selected with equal probability. If that term is present in the model at the current state of the chain then it is removed, if not then it is added. If a move is rejected then a new index is drawn, again with even probability. While this approach is computationally convenient, it is likely that for certain posterior geometries there would be significant limitations. Firstly, the proposed scheme always proposes to leave the current model. For tightly concentrated posteriors (wherein most of the probability mass is concentrated in a single model), this approach might result in low acceptance rates and inefficient exploration. A second limitation may occur in situations whereby the posterior has several modes, each corresponding to highly different model structures. In this second scenario, the proposed approach may struggle to mix from one mode to another. A potential solution would be to increase the number of indices drawn, although this could exacerbate the issues with tightly concentrated posteriors.
An alternative to the ‘bit-flipping’ approach employed here could be to incorporate the idea of neighbourhood proposals [28, 29]. These methods generalise the bit-flipping scheme above to balance exploration of the model space with exploitation near to high-probability models. These schemes typically consider moves from the current model to one in a nearby ‘neighbourhood’ that is adaptively varied based on data or posterior probabilities. These ideas are related to the more general idea of adaptive MCMC schemes [30] that seek to improve the efficiency of MCMC samplers by including adaptive proposal mechanisms.
BINDy algorithm
The overall sampling approach is summarised in Algorithm 1. In the algorithm it is assumed that there is a Gaussian prior over the parameters $\Xi$ and an inverse gamma prior over the noise variance. Note that because the parameter vector does not appear in (22), there is no need to specify an initial condition for the parameter vector.
Algorithm 1 Bayesian identification of nonlinear dynamics
@@ 3.1 A framework for Bayesian identification of nonlinear dynamics
Parameter prior mean and variance $\mu_{m}^{(0)},\Sigma_{m}^{(0)}$, noise variance prior parameters $a^{(0)},b^{(0)}$, prior over models $p(m)$, jump kernel $J(m^{\prime}|m)$, term library $\Theta(x)$, state derivatives $\dot{x}$, initial chain state ($m_{0}$, $\sigma_{0}^{2}$).
Initialise the chain at $m=m_{0}$ and $\sigma^{2}=\sigma_{0}^{2}$.
@@
     Propose a model move $m\rightarrow m^{\prime}$ by sampling from $J(m^{\prime}|m)$.
     Sample a new parameter vector $\Xi^{\prime}$ for the model $m^{\prime}$ by sampling from $p(\Xi|\mathcal{D},\sigma^{2})$ in (18).
     Accept the model and parameter move jointly with probability $\alpha(\Xi\rightarrow\Xi^{\prime})$ given by (22) else remain in place.
     Propose a new noise variance $\sigma^{2}$ using the Gibbs update given by (24).
@@
return Samples from $p(\Xi,m,\sigma^{2}|\mathcal{D})$
Posterior interpretation
Once the practitioner has collected a sufficient set of samples (convergence of the chain can be verified by observing the trace of the parameter values in $\Xi$ for example), the posterior density can be examined in the usual Monte-Carlo fashion by counting terms and by forming histograms over model parameters. In the context of nonlinear system identification, it is often of interest to compare the performance of the identified models in simulation. The simulation performance of a model is a better measure of the quality of the fit to data and as such represents a more robust challenge than prediction of state derivative from known states and library. Because the proposed approach here generates a probability distribution and not a single model, the simulation performance of the posterior can be evaluated sample-wise, by integrating samples from $p(\Xi,m,\sigma^{2}|\mathcal{D})$ forward in time (with or without added noise depending on the application). The overall simulation performance can then be established by considering the distribution over sample trajectories. Access to robust uncertainty quantification in simulation has important applications of system identification including active learning [16], model predictive control [31] and structural health monitoring [32].
@@ 3.2 Related work
The method presented herein is also not the first use of RJMCMC in the Bayesian identification of dynamic models. Troughton and Godsill [23] and Dahlin et al.  [37] consider an RJMCMC scheme for the identification of linear autoregressive models, and a RJMCMC scheme is employed in [38] to select between two candidate nonlinear models. Recently in [39], an RJMCMC approach is used to select terms in a state-space formulation using Kalman filtering techniques. The paper has the advantage that partial state observation is handled naturally however the technique is limited to linear systems only. However, to the author’s best knowledge, the present work represents the first application of RJMCMC to the SINDy-type formulation of [1], and the first to explicitly enumerate a full posterior distribution over both model terms and their parameterisation jointly in that context.
@@ 4 Results
The prior structure of the inference is set up initially to be weakly informative. In particular, the values,
$\mu_{m}\triangleq\bm{0}^{(0)},\quad\Sigma_{m}^{(0)}\triangleq 10^{3}\times I{\color[rgb]{0,0,0},}$ (29)
$a^{(0)}\triangleq 0,\quad b^{(0)}\triangleq 0{\color[rgb]{0,0,0},}$ (30)
are chosen. The prior distribution over models is selected initially to be flat with $p(m)\propto 1,\ \forall\ m\in\mathcal{M}$. The authors would note that this choice does not promote model parsimony a priori. The chain is initialised with all terms in $\Theta$ present in the model and with the noise variance set to $\sigma^{2}=1$. The Gibbs sampler is run for $6\times 10^{3}$ steps in total, with the first thousand samples discarded to remove the effects of any transient ‘burn-in’ behavior. The authors remark that this length of chain appears more than adequate in terms of convergence to the target distribution (See the appendices for a numerical study).
@@ 4.1 Legendre polynomials
The first case-study is a static sparse polynomial regression problem in which $\Theta(x)$ is static and is set to be the first 10 Legendre polynomials on the interval $x\in[-1,1]$. In order to demonstrate the recovery of small-valued parameters, a coefficient vector is selected with both zero and near-zero elements. The coefficient vector was generated randomly by the following scheme:
  • 10 random parameters were sampled uniformly on [0,1].
  • 4 randomly selected parameters (with replacement) were set to 0.
  • 2 randomly selected parameters (with replacement) were multiplied by 0.01 to drastically reduce their magnitude.
  • All parameters are rounded to 3 decimal places (so that they can be exactly reproduced in this document).
The resulting coefficient vector is given by, [^4] For clarity and readability, the results of the case study here considers only a single parameterisation of this polynomial. An additional numerical study considering many such parameterisations is presented in the appendices.
$\Xi^{*}\triangleq\left[0.549,0,0.603,0.545,0.424,0.006,0,0,0,0.004\right]^{\mkern-1.5mu\mathsf{T}}{\color[rgb]{0,0,0}.}$ (31)
@@
$y\triangleq\Theta(x)\Xi^{*}+\epsilon\quad\epsilon\sim\mathcal{N}(\bm{0},\sigma^{2}_{n}I){\color[rgb]{0,0,0},}$ (32)
@@
The performance of both E-SINDy and the proposed method are compared in Figure 1. As can be seen in the figure, Both the proposed method and E-SINDy methods are able to correctly assign low inclusion probability to the terms in $\Xi^{*}$ that are equal to zero. The posterior samples generated by BINDy are tightly concentrated around the true values such that they are virtually indistinguishable from their median. However, the STLSQ algorithm is unable to identify terms with parameters lower than the threshold parameter $\lambda$, which in this case study is set to the pysindy default value of 0.1. While this is a simplified example, it serves to demonstrate a particular pathology, which may appear when using parameter value as a proxy for term importance. However, the E-SINDy method shows multi-modality, producing some samples in clusters away from the true value. This can be attributed to the library bagging operation. Every subset of the library functions produces a sample (none are rejected - there is no Metropolisation step). This has the effect that some samples are generated with important terms missing, biasing the parameter values of the remaining terms.
@@ 4.2 Lynx-hare population dynamics
$\begin{split}\frac{dL}{dt}&=c_{11}L+c_{12}HL\\ \frac{dH}{dt}&=c_{21}H+c_{22}HL\end{split}{\color[rgb]{0,0,0},}$ (33)
where the $c_{ij}$ are unknown constants in the model. The dataset represents a significant challenge in that only 21 measurements are available.
@@ 4.3 Lorenz attractor dynamics
The third case study considers the chaotic dynamics of the Lorenz attractor. The dynamics of the Lorenz attractor are well studied in the SINDy literature [1, 16], and have become a benchmark system for advancements to the SINDy method. The dynamics are driven by the following system of ordinary differential equations:
$\begin{split}\frac{dx}{dt}&=10(x_{2}-x_{1}){\color[rgb]{0,0,0},}\\ \frac{dy}{dt}&=x_{1}(28-x_{3}){\color[rgb]{0,0,0},}\\ \frac{dz}{dt}&=x_{1}x_{2}-\frac{8}{3}x_{3}{\color[rgb]{0,0,0}.}\end{split}$ (34)
@@
For the case study in this work, the Lorenz equations above are simulated for 10 seconds (with a further 5 seconds of unseen data for extrapolation) from an initial condition of $[x_{1},x_{2},x_{3}]^{\mkern-1.5mu\mathsf{T}}=[-8,7,27]^{\mkern-1.5mu\mathsf{T}}$ at a sample rate of $10^{2}$ Hz. Overall, $10^{3}$ points are available for training. As before, $\Theta(x)$ is set as a polynomial library, considering terms up to third order. Before training, noise at 2.5% RMS is added to the state variables. The state derivatives are computed numerically using a polynomial smoothed finite-difference scheme with a difference order of 2, a window size of 5 and a polynomial order of 3. In order to reduce the error introduced by the numerical computation of the derivatives close to the edges of the signal, a few time points [^6] Dropping $12n+$ 1 points from each end for an $n$th-order finite difference scheme is the default used in pysindy [19] and here. After [16], a 2nd order finite difference scheme is used. are dropped from the signal endpoints. For the SINDy results, the threshold parameter is set to 0.2 (as is done in [16]). All other parameters are unchanged from the previous case-studies.
@@
For this third case study, the effect of the prior over models $p(m)$ is also considered. In addition to the weakly informative prior used in the previous case-study examples, a regularising geometric prior over the total number of non-zero elements in the model is also considered. The probability mass function of the regularising prior is given by,
$p(m)=(1-\theta)^{d}\theta{\color[rgb]{0,0,0},}$ (35)
@@
where $d$ is the number of terms in model $m$, and $\theta$ is a hyperparameter, set to the strongly regularising value of 0.99 in this work. The effect of this prior over models is to strongly penalise (by assigning low prior probability to) models with many terms.
@@
The result of the UQ for the Lorenz equations is compared in Figure 5 for BINDy (with and without the regularising prior) and E-SINDy. It can be seen in the figure that the unregularised BINDy assigns some posterior probability to several spurious terms, particularly in the third equation. In contrast, Both E-SINDy and the regularised BINDy assign high confidence to terms in the Lorenz equations. This is not an unexpected result. The nature of the Lorenz equations, and the use of the polynomial library results in many, highly-correlated basis functions in $\Theta(x)$. This, plus the effect of noise permit many incorrect (but low error) parametrisations of the observed dynamics. Both E-SINDy and regularised BINDy are able to overcome this difficulty by explicitly promoting model parsimony. Because the unregularised BINDy enforces no such cognitive bias, it is unsurprising that the identified posterior contains many terms. Summary statistics of the three identified posteriors are provided in Table 1.
@@
As before, sampled model trajectories (integrated forward in time) are plotted for all three methods in Figure 6. The chaotic nature of the Lorenz equations makes tracking the true dynamics highly challenging. The well-known ‘lobe-switching’ behavior of the dynamics means that small errors can quickly cause significant deviations as the dynamics become governed by a different local attractor. This effect is clearly present in the identified distributions, which quickly become multimodal after a few oscillations. Qualitatively, it is observed that the regularised BINDy samples remain in step with the true dynamics for the longest, with E-SINDy the first to diverge. This divergence can be seen almost immediately where some orange lines deviate quickly from the observed data. Indeed, this is supported by the numerical results in Table 2 which show lower error statistics for the regularised BINDy approach. In the extrapolation region, the distributions have become very diffuse, in keeping with what might be expected for uncertain forecasting of a chaotic attractor such as the Lorenz equations.
@@ 5 Discussion
As has been demonstrated in the third case-study of this work, the choice of prior over models $p(m)$ may be critical in enforcing parsimony. Here, a strongly regularising geometric distribution parameterised by the number of terms in the model has been applied successfully. However, vastly many more parametrisations are possible. It is expected that prior selection in practice will be driven by expert domain knowledge, wherever possible. For example, one could set the prior over each term directly, assigning high prior probability to terms that are expected to be in the model and low probabilities otherwise. Alternatively one could promote parsimony in some types of library functions while permitting dense models in other types. The flexible nature of the Bayesian formulation presented in this work, allows modellers to encode both domain knowledge and cognitive biases a priori. An interesting avenue for further investigation is hierarchical prior structures whereby the model-space prior is permitted to be conditional on the values its parametrisation.
@@
A related area for future investigation is the choice of the ‘jump’ kernel $J(m^{\prime}|m)$. The ‘bit-flipping’ method applied herein has proved to be effective in the case-studies considered so far, but it might be ineffective when significant multi-modality is expected in the posterior, for example when very differently parameterised models describe the data well, low probability intermediate steps might inhibit exploration of the posterior. Instead, it is interesting to imagine more sophisticated schemes that could better suit problems of this type. Also interesting are so-called ‘non-reversible’ jump methods [42] that use an adaptive scheme to ensure that the posterior is explored efficiently. Another possibility for the jump kernel is to include an approach based on the adaptive neighbourhood proposals of [28, 29]. For very large model libraries (whereby evaluation of a model with all the terms included may be computationally infeasible), such approaches may give practitioners a method to retain a computationally viable sampler, by considering model moves only to ‘nearby’ models.
The computational complexity of the proposed approach is of interest. The core computational cost in each iteration of the proposed algorithm is the matrix inversion required to perform the conjugate update of the proposed parameters in each iteration of the sampler. Because this cost is (naively) cubic $O(d^{3}+Nd^{2})$ for an RJMCMC move to a model order of $d$ (from a model order less than $d$) given $N$ observations, the worst-case analysis is that the method is $O(D^{3})$ for a library of size $D$ terms. However, because we expect many of the samples in the posterior to be sparse $(d<<D)$ the actual computational complexity (per sample) is likely to be significantly lower than this. Despite the lack of a concrete computational complexity, the authors remark that empirically the method is extremely fast. The authors are able to run the sampler at over 200 iterations per second on a single core of a Dell XPS laptop with an i7 processor.
Although the case studies in this paper have focussed on systems of first order ODEs, it is of interest to consider applications to other types of dynamical systems. Since their initial conception, SINDy-type methods have been extended to include PDEs [11], implicit dynamics [12], discrete dynamics and weak PDE solutions [13]. Because the approach here can be thought of as a replacement for the sparse regression at the core SINDy modelling there is no reason that it cannot be applied in any of these cases. However, there are some important considerations. For example in the case on implicit dynamics, care would need to be taken to treat the ill conditioning of the implicit dynamics in a robust probabilistic manner. The authors see these areas as very interesting areas for future investigation. It is also interesting to imagine the application of the proposed method to systems of very many differential equations. Such high-dimensional problems pose a challenge in that the libraries under consideration would grow combinatorially with the number of degrees of freedom. In this setting, the adaptive neighbourhood proposals in [28, 29] may encourage better mixing in the chain.